a compact car has a mass of 800kg and its efficiency is rated 18%(that is 18% of the available fuel energy is delivered to the wheels)find the energy required to accelarate the car from rest
1
Expert's answer
2020-08-06T16:45:42-0400

Let's denote the target speed with "v". Then, at this speed the kinetic energy of the car will be:


"K = \dfrac{mv^2}{2}"

where "m = 800 kg" is a mass of the car. Since its efficiency is rated "18\% = 0.18", then the energy, required to accelarate the car from rest to the speed "v" is:


"E = \dfrac{K}{0.18} = \dfrac{5mv^2}{18}[J]"

One can substitute any required speed "v" in m/s into the last expression and obtain the numerical result.

Answer. "E = \dfrac{5mv^2}{18} [J]"